I need to know how to do the standard deviation and the coefficient of variation using the data from question 3. Please show all work using formulas not financial calculators.
2. (5 points) During the early 1980s interest rates were much higher. Suppose an APR of 1 1.25% 10.00% with monthly compounding, which would you prefer (all else equal)? the effective annual rates. Bark o.. EAR, 1.75% Cwo.twK Bank One offered Show this with Two o with annual compounding on a deposit, while Bank Two offered an APR of 3. (8 points) Given the following year end stock closing prices for Mobile Tech Inc., calculate the annual holding period returns for each year from 2015 to 2018. Year 2018 2017 2016 2015 2014 HPR 474,2。 Price $45 $41 $24 $28 $20 2 5. (5 points) For the above holding period returns, what is the effective annual return if you purchased the stock in 2014 and sold it in 2018 at the prices listed? 6. (14 points) From question 4, what is the average annual HPR, the standard deviation of HP and the coefficient of variation? v Aane Irk - 31.257 Rs,
Answer #1

Answer:

Average return=Sum of HPR/4=106.30%/4=26.58%

variance- -R74

Variance=10.23%

Standard Deviation= V varance=31.98%

Year HPR X P=X-R B=P*P
2014
2015 40% 13.42% 1.80%
2016 -14.29% -40.86% 16.70%
2017 70.83% 44.26% 19.59%
2018 9.76% -16.82% 2.83%
Average return R 26.58% Variance 10.23%
Standard deviation 31.98%

Coefficient of variation= Standard deviation /Average return=31.98%/26.58%=120.34%